# From Evidence to Verdict: An Agent-Based Forensic Framework for AI-Generated Image Detection

### v1

@@ Introduction
The forensic Toolbox is a collection of specialized modules that the Evidence Gatherer can invoke to analyze the input image. The tools can be broadly categorized into the following four classes:
  • Reverse Image Search Tools: Assess image provenance by querying external online sources.
  • Metadata Extraction Tool: Parse metadata information from the image file.
  • Pre-Trained Classifiers Tool: Apply static models to assess authenticity.
  • VLM-Based Reasoning Tool: Perform semantic-level analysis using VLMs.
First, reverse image search tools are designed to identify the provenance and distribution history of input images by searching for exact matches or visually similar content across the Internet. These tools are particularly useful for determining image provenance, e.g., AI-generated images may appear on generative art platforms, while real images are more likely to be found on reputable news or photographic sites. We implement two complementary reverse image search tools. The first tool leverages the Google Cloud Vision API’s Web Detection service [7], which provides programmatic access to Google’s extensive image indexing capabilities. It primarily returns exact matches of the query image by analyzing its visual features and comparing them against Google’s web-scale image database. The API outputs structured data such as webpage titles, URLs, and contextual information where identical or highly similar images are found. However, the returned information can sometimes be sparse or insufficient for reliable judgment. To address this limitation and broaden the search scope, we implement a second approach inspired by Xu et al. [55] that simulates authentic human user behavior through web interface automation. This tool directly interacts with the Google Images search interface using Playwright, [^2] https://playwright.dev/ a modern web automation framework, to perform searches beyond the API’s restricted results. Unlike the first tool, this approach captures not only exact matches but also a wider range of visually similar images, thereby providing richer contextual evidence. It automatically uploads the input image, captures the results section from the returned webpage while filtering out advertisements and noise, and summarizes key information such as image match identification and the provenance of similar images.
Metadata analysis tool focuses on extracting and analyzing technical metadata embedded within image files to identify authenticity markers that distinguish real photographs from AI-generated content. Digital cameras and imaging devices typically embed rich metadata (EXIF data) including camera settings, GPS coordinates, timestamps, and device-specific information that are often absent or inconsistent in synthetically generated images. For example, as shown in Table 1, a genuine photograph may contain entries like EXIF:Make and EXIF:Model indicating the camera manufacturer and model. They may also include realistic optical parameters or position data such as EXIF:LensInfo and Composite:GPSPosition that are indicative of a real-world capture event. In contrast, AI-generated images frequently lack such detailed metadata or contain synthetic traces. For instance, some AI images explicitly embed generation information or even the original text prompt used for synthesis, e.g., EXIF UserComment. These anomalies serve as strong indicators of AI generation.
Our metadata extraction tool employs ExifTool, [^3] https://exiftool.org/ a comprehensive metadata manipulation library, to perform deep analysis of image files. Rather than extracting all available metadata fields, which can be overwhelming and include irrelevant information, our system implements a selective extraction strategy based on two key filtering mechanisms. First, we maintain a curated list of exact-match key fields $\mathrm{KEY\_FIELD\_EXACT}$ that have been empirically determined to provide strong authenticity signals. These include critical camera parameters such as focal length, aperture settings, ISO values, and camera manufacturer information. Second, we implement prefix-based filtering $\mathrm{KEY\_FIELD\_PREFIXES}$ to capture related metadata families. For example, all fields for camera manufacturer custom information are captured through the “MakerNotes” prefix. This dual-filtering approach ensures comprehensive coverage of relevant metadata while filtering out noise. The complete list of all key fields and prefixes is provided in the Appendix B.
Table 1: Comparison of metadata signals for distinguishing real and AI-generated images.
Metadata Field Example Value
Real Image Signal
EXIF:Make Canon
EXIF:LensInfo 4.1 123 3.5 6.4
Composite:GPSPosition 28.35 N, 81.59 W
AI Image Signal
JUMBF:Description AI Generated Image
EXIF:UserComment "A full body of a cat..."
Pre-trained classifiers tool leverages pre-trained models specifically fine-tuned for AI-generated image classification. These models have been trained on diverse datasets containing both authentic and AI-generated images, enabling them to learn discriminative features that distinguish between the two categories. We select the top five most downloaded classification models for AI-generated image detection available on Hugging Face [1]:
  • haywoodsloan/ai-image-detector-deploy [3]
  • Organika/sdxl-detector [12]
  • legekka/AI-Anime-Image-Detector-ViT [4]
  • Smogy/SMOGY-Ai-images-detector [13]
  • NYUAD-ComNets/NYUAD_AI-generated_images
    _detector
     [11]
Our implementation employs an ensemble approach that integrates multiple transformer models to enhance detection robustness and accuracy. The ensemble includes models trained on different synthetic image generation techniques, ensuring comprehensive coverage of various AI generation paradigms including GANs, diffusion models, and other neural synthesis methods. Each transformer model performs independent classification, producing probability distributions over the binary classification space. The final prediction score is obtained through a weighted voting mechanism, where each model’s contribution is scaled by its weight parameter. This weighted ensemble score is computed as:
$\text{Prediction Score}=\frac{\sum_{i=1}^{N}w_{i}\cdot s_{i}}{\sum_{i=1}^{N}w_{i}},$
@@
where $w_{i}$ represents the weight of the $i$-th model $\theta_{i}$, $s_{i}$ is the AI confidence score from $\theta_{i}$, and $N$ is the total number of loaded models. Each model is assigned an equal weight $w_{i}=1$ to ensure balanced contribution and avoid over-reliance on any individual model during inference.
VLM-based reasoning tools employs VLMs to perform sophisticated visual analysis. These models leverage their extensive training on diverse image-text pairs to identify subtle visual patterns, artifacts, and inconsistencies that may indicate AI generation. Our implementation utilizes GPT-4.1 [8] to conduct detailed image analysis. A crafted prompt guides the model to focus on specific visual characteristics that are indicative of AI generation versus authentic photography. The complete prompt is provided in Table 15 in the Appendix. When the model identifies an image as AI-generated, it is prompted to provide detailed evidence in the form of specific visual artifacts. These include unnatural textures or patterns that deviate from expected material properties, inconsistent lighting or shadow directions that violate physical principles, anatomical errors in human or animal subjects, unusual distortions in object boundaries, text rendering abnormalities, symmetry issues, and contextual inconsistencies in background elements. For images classified as real, the model is instructed to explain the supporting visual characteristics, such as realistic anatomical proportions and coherent environmental context. Unlike the other forensic tools, VLMs not only enable binary classification but also provide detailed justifications and fine-grained visual analyses that enhance the interpretability of the final decision. Additionally, the model provides a confidence assessment (high, medium, or low) based on the strength and clarity of the observed evidence. The deterministic configuration (temperature = 0, seed = 42) ensures consistent and reproducible analysis results across multiple runs.
@@
To comprehensively evaluate our agentic framework, we consider both controlled and real-world scenarios. We therefore construct our benchmark dataset to cover two distinct settings: in-the-lab and in-the-wild. The in-the-lab setting consists of images collected from well-curated, controlled datasets commonly used in prior research, while the in-the-wild setting comprises images sourced from diverse, unconstrained online platforms, reflecting the complexity and unpredictability of real-world data.
Our curated dataset comprises a total of 6000 images, evenly distributed across the two settings, with each containing 1500 AI-generated images and 1500 real images. For the in-the-lab setting, real images are sampled from well-established, curated datasets frequently used in image generation and detection research. Specifically, we sample 500 images each from Flickr30k [40], ImageNet [20], and the DIV2K dataset [16]. The AI-generated counterparts are obtained from GenImage [61] and FakeBench [33]. We sample 100 images from each of the 8 generative models included in GenImage, and 70 images from each of the 10 models in FakeBench, resulting in a total of 1500 AI-generated images.
[…]
Table 2: Dataset construction statistics.
Source Datasets / Platforms Type # Images
In-the-lab Setting
Flickr30k [40] Real Image 500
ImageNet [20] Real Image 500
DIV2K [16] Real Image 500
GenImage [61] AI Image 800
FakeBench [33] AI Image 700
In-the-wild Setting
Holopix50k [26] Real Image 500
Flickr [6] Real Image 500
Wikimedia Commons [14] Real Image 500
Lexica [9] AI Image 500
NightCafe [10] AI Image 500
Civitai [5] AI Image 500
Experimental Setup
Implementation Details. To evaluate the performance of our AIFo framework, we conduct experiments on the constructed benchmark dataset, comparing our agentic approach against a range of baseline methods. Our framework is implemented using LangGraph, [^4] https://langchain-ai.github.io/langgraph/ which is specifically designed for building stateful, multi-agent applications with LLMs. We select GPT-4.1 (gpt-4.1-2025-04-14 version) as the backbone for our LLM agents, and we set the temperature to 0 and the seed to 42 to ensure deterministic behavior across runs.
Evaluation Metrics. We formulate the evaluation as a standard binary classification task, where the goal is to assess the framework’s ability to distinguish between AI-generated and real images. Specifically, we define correctly identifying an AI-generated image as a True Positive (TP) and correctly identifying a real image as a True Negative (TN). Based on these definitions, we compute the following standard evaluation metrics: Accuracy, Precision, Recall, and F1-score. In addition to evaluating the framework’s overall performance on the entire dataset, we further analyze its behavior separately on the in-the-lab and in-the-wild subsets to assess its robustness across controlled and real-world scenarios.
Baseline Methods. To comprehensively assess the effectiveness of our proposed multi-agent framework, we compare it against a diverse set of representative baseline methods spanning both conventional and vision-language modeling paradigms. For traditional classifier-based approaches, we include CNNSpot [53], DE-FAKE [46], and PatchCraft [59], with their details discussed in Section 7.1. For VLM baselines, we adopt state-of-the-art multimodal models including GPT-4o and GPT-4.1, which have demonstrated advanced visual reasoning capabilities. During evaluation, each model is provided with the input image along with the prompt Is this a fake or real image?, and instructed to return a structured binary classification result.
Results
@@
CNNSpot [53] 0.5277 0.9826 0.0563 0.1066 0.5553 0.9882 0.1120 0.2012 0.5000 0.5000 0.0007 0.0013
PatchCraft [59] 0.6517 0.7423 0.4647 0.5715 0.8123 0.8704 0.7340 0.7964 0.4910 0.4780 0.1953 0.2773
DE-FAKE [46] 0.7142 0.6820 0.8027 0.7374 0.6720 0.6673 0.6860 0.6765 0.7563 0.6933 0.9193 0.7905
GPT-4.1 [8] 0.9416 0.9913 0.8910 0.9385 0.9332 0.9932 0.8723 0.9288 0.9500 0.9895 0.9097 0.9479
GPT-4o [39] 0.9483 0.9920 0.9038 0.9458 0.9537 0.9938 0.9130 0.9517 0.9428 0.9900 0.8947 0.9399
AIFo w/o debate (ours) 0.9635 0.9922 0.9343 0.9624 0.9730 0.9924 0.9533 0.9725 0.9540 0.9921 0.9153 0.9521
@@
Table 3 presents the comprehensive performance comparison of our multi-agent framework against five baseline methods across three evaluation settings. A detailed breakdown of their performance is provided in Table 17 in the Appendix. These results demonstrate the superior effectiveness of our proposed approach in AI-generated image detection under both controlled and real-world conditions.
Traditional Classifier Baselines. The traditional methods show varied performance patterns. CNNSpot [53] exhibits poor overall detection performance, close to random guessing. Its extremely low recall and F1-score indicate that it fails to effectively detect AI-generated images. DE-FAKE [46] achieves moderate performance, with an accuracy of 0.7142 and an F1 score of 0.7374, slightly outperforming PatchCraft [59]. These results highlight the inherent limitations of traditional classification methods in coping with rapidly evolving and continuously updated diverse AI-generated content in complex real-world scenarios.
Vision-Language Model Baselines. Both GPT-4.1 and GPT-4o demonstrate strong performance, with accuracy exceeding 0.94 and F1-scores above 0.93. However, the recall rates for both models are relatively lower compared to their precision, indicating that they still miss a significant portion of AI-generated images. This suggests challenges remain for these models in detecting certain types of AI-generated content. Our agent framework addresses this limitation through the multi-agent architecture, which combines the strengths of multiple specialized tools to achieve more reliable detection.
AIFo Framework Performance. Compared to all baseline methods, our AIFo framework consistently achieves the best performance across almost all metrics and evaluation settings. In the overall evaluation, AIFo attains an accuracy of 0.9705 and an F1-score of 0.9698, surpassing the strongest baseline GPT-4o by absolute margins of 2.22% in accuracy and 2.40% in F1-score. Moreover, this improvement is observed consistently in both in-the-lab and in-the-wild data, which span over 20 generative models, demonstrating the strong generalization capability of our framework. Through a training-free paradigm, our method is minimally affected by the evolution and iteration of generative models.
Impact of Multi-Agent Debate Mechanism. We also evaluate the impact of disabling the multi-agent debate mechanism on the performance. Notably, even without the debate mechanism, our AIFo framework still significantly outperforms GPT-4o. With the debate mechanism incorporated, our framework’s accuracy is further improved, particularly in in-the-wild setting, where it surpasses GPT-4o by 2.4%. These results demonstrate that the debate mechanism serves as an effective refinement stage, resolving potential reasoning uncertainties caused by conflicting evidence and enhancing the reliability of the final decision.
Laboratory vs. Wild Environment Analysis. A comparison between controlled (in-the-lab) and real-world (in-the-wild) environments reveals key differences. Traditional classifiers often fail to generalize: for example, PatchCraft’s F1-score drops from 0.7964 in the laboratory to 0.2773 in the wild, and CNNSpot nearly collapses entirely. This highlights the difficulty of transferring models trained on curated datasets to diverse real-world content. In contrast, our agent framework achieves consistently strong results, with F1-scores of 0.9735 in the laboratory and 0.9661 in the wild, across more than 20 generative models. These findings confirm the framework’s ability to generalize beyond controlled datasets and effectively handle the diversity of real-world AI-generated content.
Quantitative and Qualitative Analysis. To better understand where AIFo achieves improvements over GPT-4o, we analyze all samples misclassified by GPT-4o but correctly identified by AIFo. We find that AIFo correctly identifies 136 more AI-generated images compared to GPT-4o out of the total 6,000 samples. Among these corrected cases, approximately 38% are attributed to decisive reverse image search evidence, primarily from in-the-wild sources. Another 35% are resolved through metadata signals, which prove particularly effective for images containing detailed camera parameters or generation prompts embedded by community diffusion models. The remaining 27% rely on transformer-based classifier results as supporting evidence. In approximately 30% of all corrected cases, AIFo invokes the debate mechanism to resolve ambiguous or conflicting evidence, ultimately overturning the initial VLM-based assessment.
To further demonstrate the effectiveness of AIFo, we present qualitative examples in Figure 2 that show how the system integrates multiple sources of evidence and, when necessary, employs a debate mechanism to reach accurate conclusions. In the first real-image case, the Evidence Gatherer agent successfully collects evidence from four forensic tools, including image sources, camera parameters, VLM analysis, and classifier scores. The Reasoning Agent judges the evidence sufficiency and, by synthesizing all sources, correctly concludes that the image is real.
In the second AI-generated case, the realistic appearance misleads the baseline GPT-4o and the VLM tool, but other tools indicated an AI origin, creating conflicting signals. In this case, the Reasoning Agent judges that the evidence is inconsistent for a direct decision and instead triggers the debate mechanism. The pro-debate agent and con-debate agent present arguments supporting their respective stances, and the decisive argument came from the pro side, which highlighted the exact match of the image found on an AI platform. Ultimately, the Judge Agent weighs the conflicting evidence and reaches the correct conclusion that the image is AI-generated.
These cases highlight two key strengths of our framework: its ability to override misleading visual evidence using stronger provenance signals, and its capacity to reconcile conflicting evidence through structured debate. These capabilities are essential for reliable detection in both benchmarks and real-world scenarios.
Figure 2: Examples of our agent framework’s decision-making process, demonstrating diverse evidence integration across different image types and sources.
Tool Reliability and Decision Pattern Analysis. To gain deeper insights into the internal decision-making processes of our multi-agent framework, we conduct a quantitative analysis of the reliability and usage patterns of individual tools, as illustrated in Figure 3. We define a consistent judgment as a case in which the agent’s final decision is in agreement with the evidence provided by a given tool. The reliability rate of a tool is then computed as the ratio of its consistent judgments to the total number of valid (i.e., informative) pieces of evidence it produces. A higher reliability rate indicates that the agent frequently trusts and aligns with that tool’s output, suggesting its potential influence on the overall decision-making process. We also compute the coverage rate of each tool, defined as the proportion of cases in which the tool provides valid evidence relative to the total number of agent decisions.
From 3(a) and 3(b), metadata extraction and the first reverse search tool focused on exact matches achieve the highest reliability rates, meaning their evidence plays a significant role in shaping the agent’s final decisions when available. However, their lower coverage rates show they are only relevant to a subset of images. This is because in our benchmark dataset, only a small fraction of images contain rich EXIF metadata or have provenance information retrievable using the first reverse search tool. These tools thus act as high-precision, low-frequency decision anchors, exerting strong influence when available but not contributing in all cases. In contrast, the VLM analysis tool, pre-trained classifiers, and the second reverse search tool specialized in similar image retrieval achieve nearly full coverage. While the VLM’s visual reasoning remains relatively reliable, the outputs of the second reverse search tool and the classifiers are less dependable. This suggests that the agent will internally assign lower weights to these pieces of evidence. Hence, these tools serve as broad-coverage, moderate-confidence sources of evidence, providing additional context and support for the final decision but not dominating the reasoning process.
Overall, the agent appears to follow a tiered weighting strategy: prioritizing high-reliability tools when available, while leveraging high-coverage tools to maintain decision robustness.
(a) Tool reliability rates
(b) Tool coverage rates
Figure 3: Analysis of individual tool contributions to the agent framework: (a) reliability rates measuring agent trust in each tool’s evidence, and (b) coverage rates showing the proportion of decisions where each tool provides informative evidence.
Inference Efficiency and Cost Analysis. While AIFo achieves high detection accuracy, its architecture introduces additional computational cost compared to single model such as GPT-4o. To assess the efficiency of our framework, we measure the end-to-end latency and token usage for processing a single image. Table 4 summarizes the inference latency and token consumption per image for different methods. On average, AIFo requires 40.08 seconds per image, approximately 7.5 times slower than GPT-4o’s 5.31 seconds. The token usage also increases significantly, with AIFo consuming an average of 5230.86 tokens per image, compared to GPT-4o’s 715.05 tokens. The majority of the additional latency and token usage originates from the multi-round llm invocation and the execution of external tools. Despite the increased computational cost, AIFo achieves a consistent 2-3 % improvement in accuracy and over 4 % gain in recall. Meanwhile, it is important to note that AIFo operates in a training-free manner, eliminating the need for costly model retraining or fine-tuning as generative models evolve. Most importantly, unlike single model reasoning that relies solely on visual analysis, our framework integrates diverse sources of evidence, offering more verifiable interpretability that aligns with human forensic reasoning.

### v2

@@ Introduction
AIFo uses a modular toolbox of evidence providers. The forensic toolbox is a collection of specialized modules that the Evidence Gatherer Agent can invoke to analyze the input image. The tools can be broadly categorized into the following four classes.
Reverse image search. Reverse image search tools provide provenance and online distribution of input images by searching for exact or visually similar matches across the internet. We employ two complementary search tools: one based on the Google Cloud Vision API for high-precision matches, and one web-automation method inspired by Xu et al. [XML24] to capture a broader set of visually similar results. Such provenance information can be highly indicative of authenticity, e.g., AI-generated images often appear on generative art platforms, whereas real images are more likely found on news or photography websites.
Metadata Extraction. This tool extracts and analyzes embedded EXIF metadata (e.g., camera parameters, timestamps, GPS) to identify authenticity markers that differentiate real photographs from AI-generated images. We employ ExifTool for selective EXIF metadata extraction, focusing on key camera and capture-related fields while filtering irrelevant noise. The full list of key fields is provided in Appendix B.
Pre-trained classifiers. We integrate multiple publicly available transformer-based detectors as an ensemble (details in Appendix C). Each model independently predicts the probability of AI generation and outputs an AI-generation score $s_{i}$. A final prediction score is obtained via a weighted voting scheme with equal model weights:
$\text{Prediction Score}=\frac{\sum_{i=1}^{N}w_{i}\cdot s_{i}}{\sum_{i=1}^{N}w_{i}},$ (1)
@@
where $w_{i}$ represents the weight of the $i$-th model $\theta_{i}$, $s_{i}$ is the AI confidence score from $\theta_{i}$, and $N$ is the total number of loaded models. We use equal weights for simplicity and this ensemble design enhances robustness and reduces overfitting to specific generators during inference.
VLM-based analysis. A VLM-based tool conducts semantic image analysis to detect visual and contextual cues indicative of AI synthesis, examining artifacts such as lighting inconsistencies, unnatural textures, anatomical anomalies, and implausible contextual relationships. For real images, it identifies coherent structural, lighting, and contextual patterns. Beyond binary classification, the tool produces interpretable textual explanations and confidence estimates, enhancing the transparency and reliability of the judgments.
@@
CNNSpot [WWZOE20] 0.5277 0.9826 0.0563 0.1066 0.5553 0.9882 0.1120 0.2012 0.5000 0.5000 0.0007 0.0013
PatchCraft [ZXLQZ24] 0.6517 0.7423 0.4647 0.5715 0.8123 0.8704 0.7340 0.7964 0.4910 0.4780 0.1953 0.2773
DE-FAKE [SLYZ22] 0.7142 0.6820 0.8027 0.7374 0.6720 0.6673 0.6860 0.6765 0.7563 0.6933 0.9193 0.7905
GPT-4o [O23] 0.9483 0.9920 0.9038 0.9458 0.9537 0.9938 0.9130 0.9517 0.9428 0.9900 0.8947 0.9399
@@
To comprehensively evaluate our multi-agent framework under both controlled and open-world conditions, we construct a benchmark dataset covering two complementary settings: in-the-lab and in-the-wild. The in-the-lab setting consists of well-curated datasets commonly used in prior work, where image sources and generation processes are relatively controlled. In contrast, the in-the-wild setting comprises images collected from unconstrained online platforms, reflecting the diversity, noise, and uncertainty encountered in real-world forensic scenarios.
The benchmark contains 6,000 images in total, evenly split between the two settings, with 1,500 real and 1,500 AI-generated images per setting. For the in-the-lab setting, real images are sampled from Flickr30k [PWCCHL15], ImageNet [DDSLLF09], and DIV2K [AT17] (500 images each). AI-generated images are obtained from GenImage [ZCYHLLTHHW23] and FakeBench [LLWLWRL24], by sampling 100 images from each of the eight models in GenImage and 70 images from each of the ten models in FakeBench, yielding 1,500 AI-generated samples.
For the in-the-wild subset, real images are collected from publicly available online sources, including Flickr, Wikimedia Commons, and the Holopix50k dataset [HKURGOL20]. To ensure diversity, we select ten keywords: animal, building, food, indoor, landscape, person, plant, snow, sport, transportation and water. Images from Flickr and Wikimedia Commons are retrieved via keyword search and randomly sampled from the results. For Holopix50k, which lacks explicit semantic labels, we use BLIP [LLXH22] to categorize images into the same keyword set before random sampling. For each keyword and each source, we sample 50 images, yielding a balanced collection of real-world photographs. The corresponding AI-generated images are sourced from three popular generative art platforms: Lexica, NightCafe, and Civitai. Images from these platforms are generated by a wide range of text-to-image models, including DALL·E [RDNCC22], Stable Diffusion [RBLEO22], SDXL [PELBDMPR23], StyleGAN [KLAHLA20], and numerous community fine-tuned variants. We use the same set of ten keywords to retrieve AI-generated images and randomly sample 50 images per keyword from each platform. As summarized in Appendix D, the resulting benchmark spans over 20 generative models, enabling evaluation under both controlled and open-world conditions.
Experimental Setup
Implementation Details.. We evaluate AIFo under a unified decision protocol that reflects its full procedural reasoning pipeline. The framework is implemented using LangGraph, which supports stateful multi-agent workflows. All agents share the same LLM backbone (e.g., GPT-4o), ensuring consistent reasoning capacity across different agent roles. Unless otherwise specified, all agents are instantiated with temperature set to 0 and a fixed random seed (42) to reduce stochastic variation in reasoning traces. All external tool outputs, including reverse image search results and metadata extraction, were cached at evaluation time and treated as fixed inputs during benchmarking to ensure reproducibility.
Evaluation Metrics.. We evaluate performance using standard binary classification metrics, including Accuracy, Precision, Recall, and F1-score, where AI-generated images are treated as positive samples. While these metrics provide a high-level summary of detection effectiveness, they do not capture how or why a decision is reached. We therefore include targeted analyses that examine intermediate decision dynamics, such as when additional reasoning is triggered and how often the final verdict differs from the initial assessment.
Baseline Methods.. We compare AIFo against representative baseline approaches spanning both conventional classifiers and VLMs. For classifier-based methods, we include CNNSpot [WWZOE20], DE-FAKE [SLYZ22], and PatchCraft [ZXLQZ24]. For VLM baselines, we adopt the same LLM backbone as used by the agents (e.g., GPT-4o) to ensure a fair comparison. Each VLM receives the input image together with a binary classification prompt. We experiment with multiple prompt techniques for the VLM baselines such as Chain-of-Thought (CoT) [WWSBIXCLZ22] and report the best results, to provide a competitive reference for evaluating the effectiveness of our framework.
Results
Overall Performance. Table 1 reports the overall performance of AIFo and baseline approaches across three evaluation settings. The results in this table are obtained using GPT-4o as the shared LLM backbone for all agents. These metrics summarize the outcome of the complete decision process, in which a verdict may be produced either directly from initial evidence assessment or after invoking conflict-aware procedural reasoning. A detailed per-setting breakdown is provided in Appendix E. Traditional classifiers (CNNSpot [WWZOE20], DE-FAKE [SLYZ22], PatchCraft [ZXLQZ24]) perform poorly, with limited ability to generalize beyond curated datasets. GPT-4o achieves strong overall accuracy but tends to rely predominantly on visual semantics, leading to lower recall in cases where AI-generated images closely mimic real photographic patterns. Across all evaluation settings, AIFo achieves competitive or superior overall performance compared to all baselines. In the full evaluation, it attains an accuracy of 0.9705 and an F1-score of 0.9698, exceeding GPT-4o by 2.22% in accuracy and 2.40% in F1-score.
Figure 2: Examples of our agent framework’s decision-making process.
Table 2: Ablation Study of AIFo Components with GPT-4o Backbone
Method Acc Prec Rec F1
GPT-4o (direct) 0.9483 0.9920 0.9038 0.9458
GPT-4o + CoT 0.9510 0.9906 0.9107 0.9489
AIFo w/o Tools 0.9525 0.9913 0.9130 0.9505
AIFo w/o MAD 0.9635 0.9922 0.9343 0.9624
AIFo 0.9705 0.9920 0.9487 0.9698
Quantitative and Qualitative Analysis. To understand where the performance gains of AIFo originate, we analyze test samples misclassified by GPT-4o but correctly classified by AIFo. Among 6,000 samples, AIFo correctly identifies over 140 additional AI-generated images. Approximately 64% of these corrections are attributed to external evidence unavailable to single VLM inference. Roughly 34% of the corrected samples involve insufficient or conflicting signals across different evidence channels, which trigger the multi-agent debate process. In such cases, initial assessments are revised through structured evaluation of competing hypotheses, yielding a corrected final verdict by the Judge Agent. Table 2 further shows that removing either external evidence tools or the multi-agent debate mechanism consistently degrades performance compared to the full framework, confirming that both components contribute meaningfully to the overall gains.
Qualitative examples in Figure 2 further illustrate this process. In the first real-image case, multiple evidence sources provide consistent support for authenticity, allowing the framework to reach a direct verdict without invoking further conflict resolution stage. In the second case, the realistic appearance misleads the baseline GPT-4o and the VLM tool. However, conflicting provenance and classifier evidence induce the framework to defer commitment and invoke multi-agent debate. Through structured argumentation, the system identifies the inconsistency in the assessment and ultimately produces the correct AI-generated verdict.
Overall, this analysis shows that the gains of AIFo come from two complementary factors: the ability to override misleading visual cues using stronger non-visual evidence, and the capacity to revise initial judgments through explicit reasoning when evidence is insufficient or contradictory.
Tool Reliability and Decision Pattern Analysis. To better understand the internal decision-making processes of our framework, we analyze the reliability and coverage of individual forensic tools. We define reliability as the proportion of cases where a tool’s evidence aligns with the agent’s final decision, and coverage as the proportion of decisions for which a tool provides valid evidence. As shown in Figure 3, metadata extraction and the first reverse search tool achieve the highest reliability, indicating that their evidence strongly influences final decisions when available. However, their relatively low coverage indicates that such decisive evidence is only present in a subset of cases. These tools therefore function as high precision but sparse evidence sources. In contrast, the VLM-based analysis, pre-trained classifiers, and the second reverse search tool provide near-complete coverage across the dataset. Among them, VLM-based analysis contributes the most essential visual evidence and serves as the primary source of information in the majority of cases. This observation is further supported by a leave-one-out analysis shown in Figure 4, where removing the VLM tool leads to the most pronounced performance degradation, with accuracy dropping below 0.85 and recall below 0.70. By comparison, disabling metadata extraction, reverse image search, or pre-trained classifiers results in only modest performance degradation, indicating that these tools provide complementary rather than dominant signals. Instead of relying on any single tool, AIFo leverages procedural reasoning to balance these heterogeneous signals. Highly reliable but low coverage evidence is emphasized when present, while high coverage but noisier evidence provides contextual support when stronger signals are unavailable. This behavior emerges from the reasoning process itself and is not explicitly hard-coded into the system.
Inference Efficiency and Cost Analysis. The procedural reasoning framework incurs higher computational cost than VLM baselines due to multi-stage reasoning, multi-agent interaction, and external tool invocation. As summarized in Table 3, AIFo requires an average of 40.08 s and 5.2k tokens per image, corresponding to approximately $7.5\times$ and $7.3\times$ the cost of GPT-4o, respectively. While this overhead is non-negligible, it should be interpreted in the context of the framework’s adaptive decision process. As shown in earlier analyses, only a subset of samples require conflict resolution reasoning, while the majority are resolved after the initial evidence assessment. More importantly, the additional computation enables capabilities that single-shot visual reasoning cannot provide. By integrating provenance, metadata, and model-based evidence, AIFo offers verifiable decision rationales and improved reliability in ambiguous cases, which are critical in high-stakes applications. These results highlight a fundamental trade-off between efficiency and decision robustness, and suggest that our procedural reasoning is most appropriate when interpretability and reliability are prioritized over real-time constraints.
Figure 3: Analysis of individual tool contributions to the agent framework.